-
CycleTrack and SDN are teaming up to make medical school admissions more transparent and accessible! Read the announcement to learn how we’re supporting the future of CycleTrack’s free application-tracking tools .
You are using an out of date browser. It may not display this or other websites correctly.
You should upgrade or use an alternative browser.
You should upgrade or use an alternative browser.
Step I BRS error?
Started by Erek94
Get help with your application
Use all the free resources available to you from SDN: articles, guides, expert advising, forums discussions, and school research.
Is there an error in the BRS explaination or am I just missing the concept here. They say the answer is 1/64 but I got 1/16 .
I got 1/16 too. Stage IV has a 50% chance of being a carrier. And the only way to get disease expression is two carriers having a baby (1/4 chance). So I got 1/4 * 1/2 * 1/2 = 1/ 16.
Is there an error in the BRS explaination or am I just missing the concept here. They say the answer is 1/64 but I got 1/16 .
I got 1/16 too. Stage IV has a 50% chance of being a carrier. And the only way to get disease expression is two carriers having a baby (1/4 chance). So I got 1/4 * 1/2 * 1/2 = 1/ 16.
It's 1/64.
It's (1/2)*(1/2)*(1/2)*(1/2)*(1/4)
Gen II for patient's mom = 1/2
Gen III for patient's mom = 1/2
Gen II for patient's dad = 1/2
Gen III for patient's dad = 1/2
Then a child has a 1/4 chance of being homo-recessive from two hetero parents.
Both generation II and generation III are 1/2 because the people are already born and must be heterozygous or there is no chance the unborn child will be affected. In a heterozygous cross you have 1 AA, 2 Aa, 1 aa. So 2/4 are heterozygous. Since the parents are both heterozygous, their child has a 1/4 chance of being homorecessive.
Last edited:
I guess my confusion is trying to wrap my head around how gen II to III is 1/2? I woudl think its 100% because one of the parents has the disease.It's 1/64.
It's (1/2)*(1/2)*(1/2)*(1/2)*(1/4)
Gen II for patient's mom = 1/2
Gen III for patient's mom = 1/2
Gen II for patient's dad = 1/2
Gen III for patient's dad = 1/2
Then a child has a 1/4 chance of being homo-recessive from two hetero parents.
Both generation II and generation III are 1/2 because the people are already born and must be heterozygous or there is no chance the unborn child will be affected. In a heterozygous cross you have 1 AA, 2 Aa, 1 aa. So 2/4 are heterozygous. Since the parents are both heterozygous, their child has a 1/4 chance of being homorecessive.
Advertisement - Members don't see this ad
I guess my confusion is trying to wrap my head around how gen II to III is 1/2? I woudl think its 100% because one of the parents has the disease.
And the other parent is heterozygous. The only 100% chance is if both parents are homozygous recessive.
If II-2 is affected (aa) and has a child with a homozygous WT (AA) that would be:
AA x aa = 100% chance offspring is a carrier
Why is that not the case?
Oh I see what you're all saying now. I agree, 1/16 looks right.
Edit: Nope, nevermind.. back to 1/64, here's why.
Generation III is 100% carrier (both maternal and paternal sides).
For some reason the tendency to assume heterozygosity resonates with me. I remember my genetics teacher explained this, but I don't remember why or how.
So, assuming all 4 maternal grandparents are heterozygotes, you now have (1/2)*(1/2) for BOTH maternal and paternal odds and another (1/2)*(1/2) for baby
Thus, (1/2)^6.
You get (1/2)^4 if you assume that 2 grandparents are homozygous dominant and the other 2 are heterozygotes though.
Last edited:
I appreciate your time and effort.Oh I see what you're all saying now. I agree, 1/16 looks right.
Edit: Nope, nevermind.. back to 1/64, here's why.
Generation III is 100% carrier (both maternal and paternal sides).
For some reason the tendency to assume heterozygosity resonates with me. I remember my genetics teacher explained this, but I don't remember why or how.
So, assuming all 4 maternal grandparents are heterozygotes, you now have (1/2)*(1/2) for BOTH maternal and paternal odds and another (1/2)*(1/2) for baby
Thus, (1/2)^6.
You get (1/2)^4 if you assume that 2 grandparents are homozygous dominant and the other 2 are heterozygotes though.
Here is the explaination the book gives. I just dont understand how there is a 1/2 chance from II-2 to III-1. Your logic obviously got the question correct for you and I cant argue that. I would just continue to get this type of problem wrong over and over again if presented like this.